OpenStudy (anonymous):

Hamlet opened a credit card at a department store with an APR of 17.85% compounded quarterly. What is the APY on this credit card? (1 point) 35.70% 23.65% 19.08% 4.46%

OpenStudy (mathstudent55):

Here's the formula for compound interest. \( A = P \left( 1 + \dfrac{r}{n} \right) ^{nt} \) Use P = 1, r = 17.85% = 0.1785, n = 4, and t = 1. Then compare A with P.

OpenStudy (amistre64):

\[(1+\frac{apr}{n})^n=1+apy\]